OpenStudy (anonymous):

The third term of an arithmetic sequence is 21, and the eighth term is 56. The first term is _____. ***So i have no idea how to solve hahaa please help and explain! :)

OpenStudy (anonymous):

so we have two equations \[21=a+2d\] \[56=a+7d\]

OpenStudy (anonymous):

isn't 21 the 3rd term? And 56 the 8th term? how come it says 2d and 7d?

OpenStudy (anonymous):

using \[a _{n}=a+(n-1)d\] \[a _{3}=a+(3-1)d=21\] \[a _{8}=a+(8-1)d\]

OpenStudy (anonymous):

=56

OpenStudy (anonymous):

ohhhh i see :) so how do i solve for the first term?

OpenStudy (anonymous):

you can subtract (1) from (2)

OpenStudy (anonymous):

eliminating a

OpenStudy (anonymous):

i'm confused :(

OpenStudy (anonymous):

you can re-write \[a=21-2d\] and substitute in \[56=a+7d\] \[56=(21-2d)+7d\]

OpenStudy (anonymous):

and this is to solve for the 1st term? @Jonask

OpenStudy (anonymous):

the last equation is in terms of b so it solves b

OpenStudy (anonymous):

and b is the first term?

OpenStudy (anonymous):

sorry d

OpenStudy (anonymous):

oh okay and d is the first term theN??

OpenStudy (anonymous):

d is the common difference

OpenStudy (anonymous):

and when i simplify the equation i get 56=21+5d right??

OpenStudy (anonymous):

oh okay so how do i solve the first term?

OpenStudy (anonymous):

yes get d first

OpenStudy (anonymous):

okay and here is to solve for d...